find the value of
sin25+sin210+sin215......+sin2180

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Sin(theta) = sin (180- theta)

And sqr( sin(theta)) = sqr( cos (theta))

2( sqr( sin(5) + sqr(sin( 10) ...sqr(sin( 85)) + sqr(sin(90)) + sqr( sin( 180))

2( 8.5) + 1 = 18

Regards